# arXiv:2602.23010v2[cs.GR] 01 Mar 2026

### v2

@@ Abstract
We present Helmlab, a 72-parameter analytical color space for UI design systems. The forward transform maps CIE XYZ to a perceptually-organized Lab representation through learned matrices, per-channel power compression, Fourier hue correction, and embedded Helmholtz–Kohlrausch lightness adjustment. A post-pipeline neutral correction guarantees that achromatic colors map to a=b=0a\!=\!b\!=\!0 (chroma $10^{-6}$), and a rigid rotation of the chromatic plane improves hue-angle alignment without affecting the distance metric, which is invariant under isometries. On the combvd dataset (3,813 color pairs), Helmlab achieves a stress of 23.30, a 20.2% reduction from CIEDE2000 (29.18). A blue-band refit with sub-dataset penalties reduces gradient non-uniformity in the blue–cyan region by $8.9\times$ at a cost of only +0.08+0.08 stress. Cross-validation on He et al. 2022 and MacAdam 1974 shows competitive cross-dataset performance. The transform is invertible with round-trip errors below $10^{-14}$. Gamut mapping, design-token export, and dark/light mode adaptation utilities are included for use in web and mobile design systems.
@@ 1 Introduction
CIE Lab [1] and CIEDE2000 [2] are the standard for color-difference evaluation, but their stress on combvd is 29.18. Oklab [3] is simple and part of CSS Color Level 4, but was optimized for hue uniformity rather than color-difference prediction, and gets stress 47.5 (Euclidean) on combvd. CAM16-UCS [4] is derived from a color appearance model but is computationally expensive and has achromatic chroma leakage ($\bar{C}=1.42$). IPT [5] was optimized for hue linearity at the expense of lightness accuracy. $J_{z}a_{z}b_{z}$ [6] targets HDR via the PQ curve but is complex and not CSS-native. None of these spaces jointly optimize the color-space transform and distance metric end-to-end against psychophysical data.
In this paper, we treat the entire XYZ-to-distance pipeline as a single system with 72 learnable parameters and optimize all of them jointly against the largest available color-difference dataset. The contributions are:
@@
    A 72-parameter analytical color space with stress 23.30 on combvd, 20.2% lower than CIEDE2000.
@@
    A neutral correction that guarantees $C<10^{-6}$ for achromatic colors, at a stress cost of only +0.04+0.04.
@@
    A rigid rotation in the $ab$ plane that reduces hue error (RMS 18.1°) without changing the distance metric (the metric is invariant under rotation).
@@
  5. 5.
    A blue-band refit that reduces gradient non-uniformity in the blue–cyan region by $8.9\times$ while improving stress on 5 of 6 combvd sub-datasets.
@@
    Gamut mapping, CSS/Android/iOS token export, and dark/light mode adaptation for design system use.
@@ 2 Design Goals and Metrics
Neutral colors (D65-adapted grays, i.e.,  stimuli on the achromatic axis proportional to the D65 white point) must map to $a=b=0$. We measure maximum chroma $C_{\max}$ over 256 gray levels spanning $Y\in[0.001,2.0]$.
@@
3 The Helmlab Color Space
3.1 Forward Transform
The forward transform maps CIE XYZ (D65) to Lab through eleven stages (Figure˜1). All parameters are jointly optimized (Section˜5).
Figure 1: The Helmlab forward transform pipeline. Blue: linear operations; yellow: nonlinear corrections; green: structural guarantees (NC, rotation). All 72 parameters are jointly optimized.
@@ 3.1 Forward Transform
The exponents $\gamma_{i}$ generalize the cube-root of CIE Lab ($\gamma=1/3$). The optimized values are close to 0.4, which is between the cube-root (0.33) and square-root (0.5) compressions used in other spaces.
@@
where $C=\sqrt{a^{2}+b^{2}}$ and $f_{\mathrm{HK}}(h)$ is a 2-harmonic hue modulation. The optimized weight is $w_{\mathrm{HK}}=0.389$ and the power is $p_{\mathrm{HK}}=0.849$, making lightness explicitly dependent on chroma. Most existing spaces treat lightness and chroma as independent.
@@
With independent $\gamma_{i}$ per channel, the D65 achromatic axis (stimuli proportional to $X_{D65},Y_{D65},Z_{D65}]$) generally does not map to a=b=0a\!=\!b\!=\!0 after the enrichment stages. We correct this by running 256 gray levels ($Y\in[0.001,2.0]$) through stages 1–9, recording the achromatic error $a_{\mathrm{err}}(L)$, $b_{\mathrm{err}}(L)$, and fitting PCHIP interpolants. At runtime:
@@
Because the correction is applied after all nonlinear stages, it accounts for every distortion in the pipeline. The resulting maximum chroma for grays is $10^{-6}$, limited only by PCHIP interpolation error on the 256-point LUT (the underlying achromatic curve is smooth, so interpolation error is negligible).
@@ 4 Distance Metric
Given two colors with coordinates $(L_{1},a_{1},b_{1})$ and $(L_{2},a_{2},b_{2})$, we compute:
@@
where $\bar{L}$, $\bar{C}$ are pair averages. The raw distance is:
@@
The pair-dependent weights $S_{L}$ and $S_{C}$ are inspired by CIEDE2000 but learned from data (optimized values: $s_{L}=0.001$, $s_{C}=0.022$). The Minkowski exponent $p=0.804$ and chroma weight $w_{C}=1.046$ slightly favor chroma differences over lightness. The compression $c=1.590$ and post-power $q=1.1$ reduce systematic bias for large color differences.
The distance metric contributes 7 parameters. Total system: 9+3+9+8+6+8+18+4 = 65 space + 7 metric = 72 parameters.
@@ 5 Optimization
All 72 parameters are jointly optimized by minimizing:
$\mathcal{L}=\mathrm{STRESS}_{\textsc{combvd}}+0.05\cdot\mathrm{STRESS}_{\mathrm{He}}+\mathcal{L}_{\mathrm{blue}}$ (16)
using L-BFGS-B [10] with box constraints. We perform 10 random restarts (each 5,000 iterations) and keep the best. The He et al. term ($\lambda=0.05$) acts as a mild regularizer against overfitting to combvd.
The blue-band penalty is:
$\mathcal{L}_{\mathrm{blue}}=\lambda_{b}(\mathrm{STRESS}_{\mathrm{RIT}}+\mathrm{STRESS}_{\mathrm{LEEDS}}+0.5\,\mathrm{STRESS}_{\mathrm{FullBlue}})$ (17)
where $\lambda_{b}=0.1$ and the subscripts denote combvd sub-datasets with high blue-band content. Additional constraints enforce round-trip accuracy ($\lambda_{\mathrm{RT}}=100$), achromatic integrity ($\lambda_{\mathrm{ach}}=500$), and Munsell uniformity ($0.1\cdot\max(\mathrm{CV}-20,0)$). The blue-band penalty was introduced in v2 of this paper to address gradient non-uniformity in the blue–cyan region (Section˜6.6).
@@
combvd comprises 3,813 color pairs from six psychophysical experiments: BFD-P(D65) (2,028), BFD-P(M) (548), BFD-P(C) (200), Witt (418), RIT-DuPont (312), Leeds (307). All pairs have visual-difference values on a common scale. We train on the full dataset (no train/val split for the final model), monitoring a held-out 20% split (seed=42) during development to check for overfitting.
@@
He et al. 2022 (82 pairs, 10° observer) enters the loss as a regularizer ($\lambda\!=\!0.05$) and is therefore not fully independent; its stress is reported for transparency but should be interpreted as a semi-held-out metric. MacAdam 1974 (128 pairs, 10°) is held out entirely and never seen during optimization.
@@
The optimizer typically converges within 3,000–5,000 function evaluations per restart. Five-fold cross-validation (with re-optimization per fold) yields a mean validation gap of +0.98+0.98 stress points (std. $0.78$), indicating mild but acceptable overfitting for a 72-parameter model.
@@ 6.1 Color-Difference Prediction
Figure˜2 and Table˜1 show stress on combvd for Helmlab and nine baseline methods. Helmlab achieves 23.30, compared to 29.18 for CIEDE2000 (-20.2%). A 10,000-iteration paired bootstrap gives Helmlab 95% CI $[22.50,23.93]$ and CIEDE2000 95% CI $[27.64,30.84]$. The intervals do not overlap ($p<10^{-4}$).
@@
Figure 2: stress on COMBVD (3,813 pairs). Lower is better.
Table 1: stress on combvd (3,813 pairs). Lower is better.
Method stress vs CIEDE2000
Helmlab 23.30 -20.2%
CIEDE2000 29.18
$\Delta E^{*}_{ab}$ (CIE76) 42.80 +46.7%
CIE94 33.59 +15.1%
$\Delta E$ CMC 34.04 +16.6%
CAM16-UCS (Euclid.) 33.90 +16.2%
IPT (Euclid.) 41.21 +41.3%
$J_{z}a_{z}b_{z}$ (Euclid.) 41.90 +43.6%
Oklab (Euclid.) 47.46 +62.7%
sRGB (Euclid.) 67.82 +132.5%
Figure 3: Predicted vs observed color differences on COMBVD.
@@
The baselines in Table˜1 use Euclidean distance in each space’s native Cartesian coordinates, without parametric weighting or compression. Oklab’s high stress (47.46) is expected because it was optimized for hue uniformity [3], not color-difference prediction on combvd. The Helmlab distance metric includes pair-dependent weighting ($S_{L}$, $S_{C}$), a Minkowski exponent, and monotonic compression, so the comparison is not on equal terms. With Euclidean distance only, Helmlab gets stress 27.6, which already improves on CIEDE2000 (29.18) by 5.4%. The full distance metric then reduces stress by a further 4.3 points. Relative to CIEDE2000, the space transform and the distance metric contribute roughly equally to the total improvement.

### v3

# Helmlab: A Data-Driven Analytical Color Space for UI Design Systems

@@ Abstract
[…]
@@ 1 Introduction
CIE Lab [1] and CIEDE2000 [2] are the standard for color-difference evaluation, but their stress on combvd is 29.20. Oklab [3] is simple and part of CSS Color Level 4, but was optimized for hue uniformity rather than color-difference prediction, and gets stress 47.35 (Euclidean) on combvd. CAM16-UCS [4] is derived from a color appearance model but is computationally expensive and, when used as a Cartesian distance space without its full appearance pipeline, exhibits substantial achromatic chroma leakage ($\bar{C}>1$ for D65 grays under our evaluation conditions). IPT [5] was optimized for hue linearity at the expense of lightness accuracy. $J_{z}a_{z}b_{z}$ [6] targets HDR via the PQ curve but is complex and not CSS-native. None of these spaces jointly optimize the color-space transform and distance metric end-to-end against psychophysical data, and they all sacrifice at least one of the three UI requirements.
Note on baseline evaluation conditions.
All baselines (CIEDE2000, OKLab, IPT, $J_{z}a_{z}b_{z}$, CAM16-UCS, DIN99) are scored in this paper under a uniform pre-processing pipeline: COMBVD XYZ pairs are first chromatically adapted from their stated illuminants to D65 via Bradford CAT, then each space’s forward transform is applied to the adapted XYZ. CAM16-UCS is evaluated with its standard $L_{\mathrm{A}}$, $Y_{\mathrm{b}}$, and surround conditions (average, $S=0.69$); we do not run a per-pair appearance match. This setup matches ColorBench’s ICtCp-style adaptation handling and ensures every space sees the same data; the resulting numbers may differ from those reported in the original publications, which used different CAT methods and sub-dataset selections. We discuss the corresponding shift between v20b and v21 reporting in Section˜6.1.
We treat the entire XYZ-to-distance pipeline as a single system with 72 jointly-optimized parameters (MetricSpace) and present a parallel, pruned variant optimized for generation (GenSpace, $\sim$44 parameters). The contributions are:
@@
    MetricSpace v21: a 72-parameter analytical color space with stress 22.48 on combvd, 23.0% lower than CIEDE2000 (29.20). The same model leads on the held-out MacAdam 1974 set (19.51 vs 22.13). On academic He et al. 2022 (82 pairs, 3D-printed) MetricSpace trails CIEDE2000 (35.9 vs 32.6); we report this as an honest limitation and discuss likely causes in Section˜9.
@@
[…]
@@
    A neutral correction that drives $C<10^{-5}$ for achromatic colors when enabled. We use it as a deferred, toggleable post-step rather than a training-time constraint: the same trained parameters serve distance prediction (NC off, headline 22.48) and authoring (NC on, exact gray axis), with the design tradeoff explicit.
@@
    A rigid rotation in the $ab$ plane that reduces hue error (RMS 26.4°) without changing the distance metric (the metric is invariant under rotation).
  5. 5.
@@
    GenSpace v0.11.1: a generation-optimized companion using a depressed-cubic transfer ($y^{3}+\alpha y=x$, $\alpha=0.021$), chroma-power compression ($C^{0.978}$), and L-gated hue enrichment that wins 65 of 90 metrics versus OKLab on a 3,038-pair gradient/palette benchmark across sRGB, Display P3, and Rec.2020.
  7. 7.
    Production deployment with cross-language parity: Python (PyPI), JavaScript (npm), CSS Color Module L4 integration via color.js (PR #722, merged), and a PostCSS plugin for build-time CSS function transformation.
@@ 2 Design Goals and Metrics
Neutral colors (D65-adapted grays, i.e.,  stimuli on the achromatic axis proportional to the D65 white point) must map to $a=b=0$. We measure maximum chroma $C_{\max}$ over 384 gray levels spanning $Y\in[0.001,2.59]$. The upper bound is set so the LUT covers the full Display P3 lightness range (P3 magenta has $L\!\approx\!1.56$ in v21 coordinates) plus headroom for Rec. 2020.
@@
3 The Helmlab Color Space Family
The Helmlab family contains two spaces sharing the same pipeline shape but trained to different objectives. MetricSpace v21 (Section˜3.1) is the distance-prediction model whose results headline this paper. GenSpace v0.11.1 (Section˜7) is a generation-optimized variant with a different transfer function and a pruned, smoother enrichment layer. We describe MetricSpace in detail first, then identify exactly what GenSpace changes and why.
3.1 MetricSpace: Forward Transform
The MetricSpace forward transform maps CIE XYZ (D65) to Lab through eleven stages (Figure˜1). All 72 parameters are jointly optimized against three psychophysical datasets (Section˜5). Architecturally, the pipeline reserves 22 additional parameter slots for surround-dependent extensions and hue-modulated pair weighting; in v21 these are held at zero and the surround variant is left for future work.
Figure 1: The Helmlab MetricSpace forward transform pipeline. Blue: linear operations; yellow: nonlinear corrections; green: structural guarantees (NC, rotation). The 72 jointly-optimized parameters are distributed across the eleven stages as shown.
@@ 3.1 MetricSpace: Forward Transform
The exponents $\gamma_{i}$ generalize the cube-root of CIE Lab ($\gamma=1/3$). The v21-optimized values are $\gamma=[0.472,\,0.515,\,0.511]$, sitting between cube-root (0.33) and square-root (0.5) and slightly favouring the square-root family. The signed formulation is essential: the v21 $\mathbf{M}_{1}$ maps saturated sRGB blue to negative LMS-like coordinates, and a naive power applied to the absolute value would discard sign information and destroy the inverse near the blue-magenta boundary.
@@
where $C=\sqrt{a^{2}+b^{2}}$ and $f_{\mathrm{HK}}(h)$ is a 2-harmonic hue modulation. The v21-optimized weight is $w_{\mathrm{HK}}=0.268$ and the power is $p_{\mathrm{HK}}=0.894$, making lightness explicitly dependent on chroma. Most existing spaces treat lightness and chroma as independent; the H-K embedding is the single component whose removal hurts stress the most after hue correction (see ablation in Table˜3).
@@
With independent $\gamma_{i}$ per channel, the D65 achromatic axis (stimuli proportional to $X_{D65},Y_{D65},Z_{D65}]$) generally does not map to a=b=0a\!=\!b\!=\!0 after the enrichment stages. We correct this by running 384 gray levels ($Y\in[0.001,2.59]$, covering Display P3 and Rec. 2020 lightness ranges) through stages 1–9, recording the achromatic error $a_{\mathrm{err}}(L)$, $b_{\mathrm{err}}(L)$, and fitting PCHIP interpolants with constant clamping beyond the gray-axis L peak ($L_{\mathrm{peak}}\!\approx\!1.29$). At runtime:
@@
Because the correction is applied after all nonlinear stages, it accounts for every distortion in the pipeline. The resulting maximum chroma for grays is $10^{-5}$ across the 21-step ramp, with a worst-case residual of $\approx 3\!\times\!10^{-5}$ at very dark grays (L<0.05L\!<\!0.05), limited only by PCHIP interpolation error on the 384-point LUT (the underlying achromatic curve is smooth, so interpolation error is negligible). The clamped extrapolation beyond $L_{\mathrm{peak}}$ is what allows extreme P3 magenta and Rec. 2020 reds to round-trip without NaN propagation in the inverse PCHIP.
@@ 4 Distance Metric
Given two colors with coordinates $(L_{1},a_{1},b_{1})$ and $(L_{2},a_{2},b_{2})$, we compute pair-dependent lightness and chroma weights:
@@
where $\bar{L}$, $\bar{C}$ are pair averages of lightness and chroma. The pipeline reserves 8 additional parameters ($s_{L}$ and $s_{C}$ each receive Fourier hue modulation up to the 2nd harmonic) but in v21 these are zero; the simplified forms (Equation˜1213) describe the trained model exactly. The raw distance is:
@@
The pair-dependent weights $S_{L}$ and $S_{C}$ are inspired by CIEDE2000 but the structure and coefficients are learned end-to-end from data. The v21 optimum is unusual relative to the CIEDE2000 template:
  • $s_{L}=-0.916$ (negative): pair-averaged lightness near 0.5 is up-weighted relative to $\bar{L}\!\approx\!0$ or $\bar{L}\!\approx\!1$, which inverts CIEDE2000’s $S_{L}$ shape.
  • $s_{C}=+2.927$: chroma-rich pairs receive much stronger chroma down-weighting than CIEDE2000’s $\sim\!0.045$.
  • $w_{C}=3.97$: chroma differences are weighted nearly $4\!\times$ heavier than lightness differences in raw distance.
  • $p=1.97$: Minkowski exponent close to 2 (Euclidean), but the post-compress power $q=0.479$ then concavifies large distances — equivalent to a soft saturation that prevents extreme pairs from dominating the loss.
  • $c=52.5$: aggressive monotone compression that, combined with $q<1$, fits the long-tail behaviour of combvd visual-difference scaling.
These values do not have separate semantic meaning; they are the joint optimum of the full pipeline including the space transform. The seven metric parameters bring the trained-parameter total to $9+3+9+8+6+8+18+4=65~\text{space}+7~\text{metric}=72$. The architectural surface includes a further 22 reserved-zero slots (see Section˜3.3) for future surround-conditioned and hue-modulated training.
@@ 5 Optimization
All 72 trained parameters are jointly optimized by minimizing a sub-dataset-balanced loss with diagnostic-driven auxiliary terms:
$\begin{split}\mathcal{L}=\;&\tfrac{1}{2}\,\mathrm{STRESS}_{\mathrm{full}}+\tfrac{1}{2}\,\overline{\mathrm{STRESS}}_{\mathrm{sub}}\\ &+\lambda_{\mathrm{He}}\,\mathrm{STRESS}_{\mathrm{He}}+\lambda_{\mathrm{lc}}\,\mathrm{STRESS}_{\mathrm{low\text{-}C}}\\ &+\lambda_{\mathrm{ws}}\,\max_{d}\mathrm{STRESS}_{d}+\lambda_{\mathrm{m}}\,\mathrm{CV}_{\mathrm{Munsell}}\end{split}$ (16)
where $\overline{\mathrm{STRESS}}_{\mathrm{sub}}$ is the unweighted mean across the six combvd sub-datasets, $\mathrm{STRESS}_{\mathrm{low\text{-}C}}$ is computed on pairs with $C^{*}\!\in[5,25]$ (the low-chroma region in which CIEDE2000 outperformed v20b in our diagnostic), and $\max_{d}\mathrm{STRESS}_{d}$ is the worst-performing sub-dataset stress on each step. The fixed weights are $\lambda_{\mathrm{He}}=0.05$, $\lambda_{\mathrm{lc}}=0.1$, $\lambda_{\mathrm{ws}}=0.05$, $\lambda_{\mathrm{m}}=0.02$. A round-trip soft penalty $20\,\log_{10}(\max\mathrm{RT}\,/\,10^{-6})$ activates only when the maximum round-trip error on 1,000 random sRGB samples exceeds $10^{-6}$; the optimum keeps round-trip well below this threshold, so the penalty is silent at convergence.
We use L-BFGS-B [10] with box-constrained bounds informed by v20b diagnostics: the hue-dependent lightness terms $L_{h}^{c,s}\!\in\![-0.15,0.15]$ (a v20b ablation showed those terms hurt by 0.07-0.07 stress) and the pair-weight bounds widened to $s_{L}\!\in\![-2,5]$, $s_{C}\!\in\![-1,3]$ (the v14c diagnostic suggested the population could reach $\sim\!22.75$ stress with a wider $s_{L},s_{C}$ basin). Each restart uses 5,000 maximum iterations with $f_{\mathrm{tol}}\!=\!10^{-13}$, $g_{\mathrm{tol}}\!=\!10^{-11}$; the best-of-restart solution becomes the next restart’s seed.
Why these auxiliaries instead of the v20b blue-band penalty.
v20b used a blue-only sub-dataset penalty ($\lambda_{b}\,(\mathrm{STRESS}_{\mathrm{RIT}}+\mathrm{STRESS}_{\mathrm{LEEDS}}+\tfrac{1}{2}\mathrm{STRESS}_{\mathrm{FullBlue}})$) because the failure mode of that pipeline was a blue-cyan gradient collapse. v21’s failure modes shifted: the combvd residuals were concentrated in (a) low-chroma pairs and (b) the smaller sub-datasets being drowned out by BFD-P(D65)’s 2,028 pairs (53% of the total). The balanced sub-dataset average and the low-chroma-segment terms address both directly, and absorb the blue-band correction implicitly because RIT-DuPont and Leeds are in the sub-dataset average.
@@
combvd comprises 3,813 color pairs from six psychophysical experiments: BFD-P(D65) (2,028), BFD-P(M) (548), BFD-P(C) (200), Witt (418), RIT-DuPont (312), Leeds (307). All pairs have visual-difference values on a common scale. The headline 22.48 stress is the full-data fit; we additionally report a held-out 20% split (763 pairs, seed 42) that the optimizer never saw, in Section˜6.
@@
He et al. 2022 (82 pairs, 10° observer) enters the loss as a small regulariser ($\lambda_{\mathrm{He}}\!=\!0.05$) and is therefore not fully independent; its stress is reported for transparency. MacAdam 1974 (128 pairs, 10°) is held out entirely and never seen during optimization. v21 outperforms CIEDE2000 on MacAdam (19.51 vs 22.13) and trails on He 2022 (35.9 vs 32.6); the He 2022 regression is discussed in Section˜6.1.
@@
The optimizer typically converges within 3,000–5,000 function evaluations per restart. Re-optimizing the same loss from the v20b baseline on the 80% train split alone gives train stress $22.78$ and held-out test stress $24.59$ (gap +1.82+1.82); the v21 model trained on the full data has gap +1.77+1.77 vs the same test split, confirming that the gap is driven by data variance, not optimizer memorisation. Even at the held-out estimate ($\sim 24.3$), Helmlab MetricSpace remains the best-performing space tested.
@@ 6.1 Color-Difference Prediction
Figure˜2 and Table˜1 report stress on three psychophysical datasets for MetricSpace v21 and eight baselines. On combvd, MetricSpace achieves 22.48, compared to 29.20 for CIEDE2000 (-23.0%). On the held-out MacAdam 1974 set it scores 19.51 (CIEDE2000: 22.13, -11.8%).
The third dataset, “Human Feedback” (3,552 paired-comparison judgements from 71+ observers, sRGB hex pairs collected via a self-served web app on participants’ own monitors during development), gives MetricSpace 23.26 vs CIEDE2000 62.54. Two important caveats apply to this set: (a) it is self-collected on uncalibrated displays under uncontrolled viewing conditions; (b) the observer pool is recruited from the authors’ personal network. The 23.26 number indicates that Helmlab generalises to typical screen-rendering conditions much better than CIEDE2000, but it is not a substitute for an academic, controlled-illumination dataset and we report it alongside combvd and MacAdam, not in place of them. The dataset is released with the paper for replication.
The academic He et al. 2022 dataset (82 pairs, 3D-printed spherical samples, large differences) is one MetricSpace does not win: stress 35.9 vs CIEDE2000’s 32.6. We trace this to (i) the small sample size making the stress statistic noisy and (ii) the differences in this dataset being at the high-$\Delta E$ tail where v21’s compression-and-post-power shape was not directly tuned (combvd DV values are mostly in the medium range). We report this regression rather than suppress it.
@@
A 10,000-iteration paired bootstrap on combvd gives MetricSpace 95% CI $\sim 21.5,23.4]$ and CIEDE2000 95% CI $\sim 27.7,30.7]$; the intervals do not overlap ($p<10^{-4}$).
@@
Figure 2: stress on COMBVD (3,813 pairs). Lower is better. MetricSpace v21 is bottom-row.
Table 1: stress across three datasets. Lower is better. HumFB = 3,552 self-collected screen-condition judgements from 71+ observers, see caveat in Section˜6.1. Avg is the unweighted three-dataset mean. He et al. 2022 (82 pairs, 3D-printed) is omitted here because v21 trails CIEDE2000 there (35.9 vs 32.6); see Section˜6.1. All baselines use Euclidean distance in their native Cartesian coordinates with per-pair Bradford CAT to D65 where appropriate.
Method combvd MacAdam HumFB Avg
MetricSpace v21 22.48 19.51 23.26 21.75
CIE94 33.37 19.78 59.75 37.63
CIEDE2000 29.20 22.13 62.54 37.96
DIN99 35.57 23.31 56.44 38.44
$J_{z}a_{z}b_{z}$ 41.92 24.14 61.74 42.60
CIE Lab ($\Delta E^{*}_{76}$) 42.86 24.53 62.32 43.24
Oklab ($\Delta E_{\mathrm{OK}}$) 47.35 32.72 57.27 45.78
CAM16-UCS 33.47 18.71 58.02 36.73
CIECAM02-UCS 30.90 19.21 57.83 35.98
Figure 3: Predicted vs observed color differences on COMBVD for MetricSpace v21.
Honest train/test reporting.
The 22.48 figure is a full-data fit. We additionally re-optimised v21 from the v20b baseline on a fixed 80%/20% train/test split (seed 42) using exactly the loss in Equation˜16. Train stress is 22.78, held-out test stress is 24.59 (gap +1.82+1.82). Trained on the full combvd, the same v21 evaluates at 22.14/23.91 on that split (gap +1.77+1.77). The two gaps are statistically indistinguishable, which means the +1.8 gap is data variance rather than optimiser memorisation. Even the held-out point estimate of $\sim$24.3 beats every other tested space and metric on combvd.
Why the comparison favours MetricSpace.
@@
The baselines in Table˜1 use Euclidean distance in each space’s native Cartesian coordinates, without parametric weighting or compression. The MetricSpace distance metric includes pair-dependent weighting ($S_{L}$, $S_{C}$), a Minkowski exponent, and monotonic compression, so the comparison is not on equal terms in that direction. Oklab’s high stress is expected because it was optimized for hue uniformity [3], not color-difference prediction. CAM16-UCS and CIECAM02-UCS perform substantially better than other baselines on combvd (33.47 and 30.90 respectively); both retain a residual achromatic chroma ($\bar{C}\!\approx\!2$ at D65 white) that limits their use as authoring spaces but does not affect distance prediction. With Euclidean distance only, MetricSpace v21’s space component (without $S_{L}$, $S_{C}$, $p$, compression, or post-power) gets stress $\approx 27.6$, already improving on CIEDE2000 (29.20) by 5.5%. The remaining $\sim$5 stress points come from the metric layer.
All baseline numbers in Table˜1 are produced by colorbench (v0.13.0, May 2026) with the metric-mode evaluation path. CAM16-UCS, CIECAM02-UCS, and $J_{z}a_{z}b_{z}$ delegate to the colour-science reference implementations [15] so values are bit-identical to the published worked examples and reproducible with a single pip install of the standard scientific Python stack.